OpenStudy (anonymous):

How fast must you drive to cover 700 miles in 13 hours? Round your answer to the nearest whole number.

OpenStudy (whynot):

S=VT V=S/T V=700/13 v=54miles/hour

OpenStudy (anonymous):

Thanks